# On the asymptotics of orthogonal polynomials on multiple intervals with non-analytic weights

### v2

@@ 3.1. Chebyshev-like polynomials on multiple intervals
So, suppose $\alpha_{j},\beta_{j}\in\{1/2,-1/2\}$. In the case where $f_{j}=1/h_{j}$ is analytic in a neighborhood of $[a_{j},b_{j}]$ for every $j$, local parametrices are not needed [23] (see also [14]). To treat the case of non-analytic $f_{j}$ using this fact, one needs to find an extension of $f_{j}$ to an entire neighborhood of $[a_{j},b_{j}]$ to appropriately pose a hybrid $\overline{\partial}$-RH problem. This was accomplished in [37]. While we follow this idea, we use a slightly different method of extension, since we have already developed our extension operator . First, we recall that for $\epsilon>0$, $f_{j}\in W^{k,p}([a_{j},b_{j}])$ has a $W^{k,p}([a_{j}-2\epsilon,b_{j}+2\epsilon])$ extension. Similarly, $f_{j}\in C^{k,\alpha}([a_{j},b_{j}])$ has a $C^{k,\alpha}([a_{j}-2\epsilon,b_{j}+2\epsilon])$ extension. This can be constructed simply by polynomial interpolation of the function and its first $k$ derivatives at the endpoints — the Taylor polynomial. For what follows we will assume $f_{j}\in C^{k,\alpha}([a_{j},b_{j}])$, $k\geq 1$ and any $\alpha>0$ or $f_{j}\in W^{k,q}([a_{j},b_{j}])$, $k\geq 1$ and $q>4$. Then, we find a sequence of polynomials $(p^{\prime}_{m,j})_{m\geq 1}$ of degree $m$ such that
$E_{j,m,q}:=\|f_{j}^{\prime}-p^{\prime}_{m,j}\|_{L^{q}([a_{j}-2\epsilon,b_{j}+2\epsilon])}\to 0,\quad q>4.$
From Jackson’s theorem, see [2, Theorem 3.7.2], we can select the sequence such that $E_{j,m,\infty}=\mathrm{O}(m^{k-1+\alpha})$ if $f\in C^{k,\alpha}([a_{j},b_{j}])$, but more refined estimates are possible for $W^{k,q}([a_{j},b_{j}])$ [17]. Set, for example, $p_{m,j}=f_{j}(a_{j})+\int_{a_{j}}^{x}p_{m,j}^{\prime}(x^{\prime})\mathop{}\!\mathrm{d}x^{\prime}$. Then, define
@@
$\mathbf{Z}_{n}(x,y)=\mathbf{Y}_{n}(z;\mu)\begin{cases}\begin{bmatrix}1&0\\ -f_{j}^{\ddagger}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Omega_{j,+}^{\prime},\vspace{3pt}\\ \vspace{3pt}\begin{bmatrix}1&0\\ f_{j}^{\ddagger}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Omega_{j,-}^{\prime},\\ \mathbf{I}&\text{otherwise}.\end{cases}$
@@
¯$\overline{\partial}\mathbf{Z}_{n}(x,y)=\begin{cases}\mathbf{Z}_{n}(z;\mu)\begin{bmatrix}0&0\\ -\overline{\partial}r_{m,j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}&z\in\Omega_{j,+}^{\prime},\vspace{3pt}\\ \vspace{3pt}\mathbf{Z}_{n}(z;\mu)\begin{bmatrix}0&0\\ \overline{\partial}r_{m,j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}&z\in\Omega_{j,-}^{\prime},\\ \mathbf{0}&z\in\mathbb{C}\setminus\overline{\Omega}_{j}^{\prime}.\end{cases}$
@@
$\Sigma_{j,+}^{\prime}$ =$=\{z~{}:~{}x=a_{j}-\epsilon/2,~{}0\leq y\leq\epsilon/2\}$
@@
i.e. the “top” boundary of the box $\Omega_{j}^{\prime}$. We give this negative orientation (as the boundary of $\Omega_{j}^{\prime})$. And set $\Sigma_{j,-}=\{z~{}:~{}\bar{z}\in\Sigma_{j,+}\}$, inheriting orientation.
@@
$\mathbf{Z}_{n}^{+}(x,y)=\mathbf{Z}_{n}^{-}(x,y)\begin{cases}\begin{bmatrix}1&0\\ f_{j}^{\ddagger}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Sigma_{j,-}^{\prime}\cup\Sigma_{j,+}^{\prime},\vspace{3pt}\\ \begin{bmatrix}0&\rho(x)\\ -1/\rho(x)&0\end{bmatrix}&z\in[a_{j},b_{j}].\end{cases}$
@@
We find that $\mathbf{T}_{n}(x,y)$ solves the normalized hybrid $\overline{\partial}$ problem
$\mathbf{T}_{n}^{+}(x,y)$ $=\mathbf{T}_{n}^{-}(x,y)\begin{cases}\begin{bmatrix}1&0\\ f_{j}^{\ddagger}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)}&1\end{bmatrix}&z\in\Sigma_{j,-}^{\prime}\cup\Sigma_{j,+}^{\prime},\vspace{3pt}\\ \begin{bmatrix}0&\rho(x)\\ -1/\rho(x)&0\end{bmatrix}&z\in(a_{j},b_{j}),\vspace{3pt}\\ \,\mathrm{e}^{-\Delta_{j}{\boldsymbol{\sigma}_{3}}}&z\in(b_{j},a_{j+1}),\end{cases}$
¯$\overline{\partial}\mathbf{T}_{n}(x,y)$ $=\mathbf{T}_{n}(z)\begin{cases}\begin{bmatrix}0&0\\ -\overline{\partial}r_{m,j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)}&0\end{bmatrix}&z\in\Omega_{j,+}^{\prime},\vspace{3pt}\\ \vspace{3pt}\begin{bmatrix}0&0\\ \overline{\partial}r_{m,j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)}&0\end{bmatrix}&z\in\Omega_{j,-}^{\prime},\\ \mathbf{0}&z\in\mathbb{C}\setminus\overline{\Omega_{j}^{\prime}},\end{cases}$
@@
$\mathbf{R}_{n}^{+}(x,y)=\mathbf{R}_{n}^{-}(x,y)\mathbf{G}_{n}(z)\begin{bmatrix}1&0\\ f_{j}^{\ddagger}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)}&1\end{bmatrix}\mathbf{G}_{n}(z)^{-1},\quad z\in\Sigma_{j,-}^{\prime}\cup\Sigma_{j,+}^{\prime},$
¯$\overline{\partial}\mathbf{R}_{n}(x,y)$
$=\mathbf{R}_{n}(z)\begin{cases}\mathbf{G}_{n}(z)\,\mathrm{e}^{-{\boldsymbol{\sigma}_{3}}G(z)}\begin{bmatrix}0&0\\ -\overline{\partial}r_{m,j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)-2G(z)}&0\end{bmatrix}\,\mathrm{e}^{{\boldsymbol{\sigma}_{3}}G(z)}\mathbf{G}_{n}(z)^{-1}&z\in\Omega_{j,+}^{\prime},\vspace{3pt}\\ \vspace{3pt}\mathbf{G}_{n}(z)\,\mathrm{e}^{-{\boldsymbol{\sigma}_{3}}G(z)}\begin{bmatrix}0&0\\ \overline{\partial}r_{m,j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)-2G(z)}&0\end{bmatrix}\,\mathrm{e}^{{\boldsymbol{\sigma}_{3}}G(z)}\mathbf{G}_{n}(z)^{-1}&z\in\Omega_{j,-}^{\prime},\\ \mathbf{0}&z\in\mathbb{C}\setminus\overline{\Omega_{j}^{\prime}},\end{cases}$
@@
Importantly, we note that by Lemma C.3, $\,\mathrm{e}^{-2G(z)}/\omega_{j}(z)$ is bounded in a neighborhood of $[a_{j},b_{j}]$. From the (at most) quarter-root singularities of $\mathbf{G}_{n}(z)\,\mathrm{e}^{-{\boldsymbol{\sigma}_{3}}G(z)}$ we find for $z\in\Omega_{j}^{\prime}$ that
$\mathbf{X}_{n}(x,y)=\mathbf{G}_{n}(z)\,\mathrm{e}^{-{\boldsymbol{\sigma}_{3}}G(z)}\begin{bmatrix}0&0\\ -\overline{\partial}r_{m,j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)-2G(z)}&0\end{bmatrix}\,\mathrm{e}^{{\boldsymbol{\sigma}_{3}}G(z)}\mathbf{G}_{n}(z)^{-1}$
satisfies
$\|\mathbf{X}_{n}(x,y)\|\leq C|\overline{\partial}r_{m,j}^{\sharp}(x,y)||z-a_{j}|^{-1/2}|z-b_{j}|^{-1/2}\,\mathrm{e}^{-2n\operatorname{Re}g(z)}.$
Note that, by (E.1), we have
(3.1) $\|\mathbf{X}_{n}\|_{L^{p}(\Omega_{j}^{\prime})}\leq C\|\overline{\partial}r_{m,j}^{\sharp}\|_{L^{q}(\Omega_{j}^{\prime})}\|h_{j,n}\|_{L^{r}(\Omega_{j}^{\prime})},\quad h_{j,n}(z):=|z-b_{j}|^{-1/2}\,\mathrm{e}^{-2n\operatorname{Re}g(z)},$
for $1\leq p\leq 4$, and $1/p=1/q+1/r$. Since we will need $p>2$ below, and using (E.1) requires $\gamma<2$, we choose $r<4$, $q>4$, giving
$\|\mathbf{X}_{n}\|_{L^{p}(\Omega_{j}^{\prime})}=O(n^{-1}E_{j,m,q}).$
Note that by choosing $p$ arbitrarily close to $2$, we can choose $q$ arbitrarily close to $4$.
@@
¯$\overline{\partial}\mathbf{C}_{n}(x,y)$ $=\mathbf{C}_{n}(z)\mathbf{X}_{n}(x,y),\quad z\in\Omega_{j}^{\prime}$
@@
From (2.9) we find that the operator
$\mathcal{L}_{n}\mathbf{U}:=\mathcal{K}(\mathbf{U}\mathbf{X}_{n})$
satisfies
$\|\mathcal{L}_{n}\mathbf{U}\|_{L^{\infty}(\mathbb{C})}\leq C\|\mathbf{U}\|_{L^{\infty}(\Omega^{\prime})}n^{-1}\max_{j}E_{j,m,q}.$
Therefore, if $m=\mathrm{O}(n)$, for $n$ sufficiently large there exists a unique solution $\mathbf{V}_{n}\in L^{\infty}(\Omega_{j}^{\prime})$ to the above equation that satisfies
$\|\mathbf{V}_{n}\|_{L^{\infty}(\mathbb{C})}=\mathrm{O}(n^{-1}\max_{j}E_{j,m,q}).$
Again, this implies that $\|\mathbf{C}_{n}(x,y)\|$ and $\|\mathbf{C}_{n}(x,y)^{-1}\|$ are uniformly bounded.
Next, we make a similar observation to what was made in [37]. The Bernstein–Walsh inequality, see [29, Theorem III.2.1], in its simplest form, states that if a polynomial $p(x)$, of degree $n$, satisfies
@@
Thus, for the sequence of polynomials $p_{m,j}(x)$ that converge uniformly on $[a_{j},b_{j}]$, we have $|p_{m,j}(x)|\leq C$ for some constant, independent of $m$ (and $j$). From this, and Proposition A.1, provided that $\epsilon$ is sufficiently small we have that
@@
For $k\geq 1$, $q>4$ and [^4] Recall that $\epsilon$ appears in the definition of $\Sigma_{j,\pm}^{\prime}$. $\epsilon$ sufficiently small set
$\mathbf{F}^{\natural}(x,y)$ =$=\begin{cases}\begin{bmatrix}1&0\\ \pm f_{j}^{\ddagger}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Omega_{j,\pm}^{\prime},\\ \mathbf{I}&\text{otherwise}.\end{cases}$
Then for $n$ sufficiently large,
@@
$\mathbf{V}_{n}(x,y)=\mathrm{O}\left(\frac{n^{-1}\max_{j}E_{j,m,q}}{1+|z|}\right),\quad z\in\mathbb{C},\quad m=\lfloor cn\rfloor.$
Remark 3.2.
In comparing our results, in the case $g=0,P=0$, with that in [37], we see that our error term is smaller only by a factor of $\log n$ if we use $q=\infty$ and select $h_{j}\in C^{k,\alpha}([a_{j},b_{j}])$. But, in general, we do only require our perturbations to be once differentiable almost everywhere, rather than three times.
Using bounds on $L^{p}$ polynomial approximation from, for example, [17], if $h_{j}(x)=1+\mathbbm{1}_{[a_{j},d_{j}]}(d_{j}-x)^{1+\alpha}$, one can show that if $\alpha>3/4$ then a polynomial approximation can be selected such that
$E_{j,m,q}=O(m^{-1}).$
This gives error terms of $\mathrm{O}(n^{-2})$ for this particular $h_{j}\in C^{1,\alpha}([a_{j},b_{j}])$. If we only used $h_{j}\in C^{1,\alpha}([a_{j},b_{j}])$ and $p=\infty$ the estimate would be $\mathrm{O}(n^{-1-\alpha})$. The bound of $\mathrm{O}(n^{-2})$ still does not quite close the gap seen in our numerical experiments below (see Figures 5 and 4), which indicate an error term of $\mathrm{O}(n^{-2-\alpha})$.

### v4

@@ 3.1. Chebyshev-like polynomials on multiple intervals
So, suppose $\alpha_{j},\beta_{j}\in\{1/2,-1/2\}$. In the case where $f_{j}=1/h_{j}$ is analytic in a neighborhood of $a_{j},b_{j}]$ for every $j$, local parametrices are not needed [26] (see also [17]). To treat the case of non-analytic $f_{j}$ using this fact, one needs to find an extension of $f_{j}$ to an entire neighborhood of $a_{j},b_{j}]$ to appropriately pose a hybrid $\overline{\partial}$-RH problem. This was accomplished in [40]. While we follow this idea, we use a slightly different method of extension, since we have already developed our extension operator . First, we recall that for $\epsilon>0$, $f_{j}\in W^{k,p}([a_{j},b_{j}])$ has a $W^{k,p}([a_{j}-2\epsilon,b_{j}+2\epsilon])$ extension. Similarly, $f_{j}\in C^{k,\alpha}([a_{j},b_{j}])$ has a $C^{k,\alpha}([a_{j}-2\epsilon,b_{j}+2\epsilon])$ extension. This can be constructed simply by polynomial interpolation of the function and its first $k$ derivatives at the endpoints — the Taylor polynomial. For what follows we will assume $f_{j}\in C^{k,\alpha}([a_{j},b_{j}])$, $k\geq 1$ and any $\alpha>0$ or $f_{j}\in W^{k,q}([a_{j},b_{j}])$, $k\geq 1$ and $q>4$. Then, we find a sequence of polynomials $(p^{\prime}_{m,j})_{m\geq 1}$ of degree $m$ such that
$E_{j,m,q}:=\|f_{j}^{\prime}-p^{\prime}_{m,j}\|_{L^{q}([a_{j}-2\epsilon,b_{j}+2\epsilon])}\to 0,\quad q>6.$
From Jackson’s theorem, see [4, Theorem 3.7.2], we can select the sequence such that $E_{j,m,\infty}=\mathrm{O}(m^{k-1+\alpha})$ if $f\in C^{k,\alpha}([a_{j},b_{j}])$, but more refined estimates are possible for $W^{k,q}([a_{j},b_{j}])$ [20, 10]. Specifically, if $f_{j}\in W^{k,p}([a_{j},b_{j}])$ we can select the sequence such that [10, (5.4.16)]
$E_{j,m,q}\leq Cm^{-k+1}\sum_{\ell=\min\{k-1,m+1\}}^{k-1}\|f_{j}^{(\ell+1)}\|_{L^{q}([a_{j}-2\epsilon,b_{j}+2\epsilon])}\leq C^{\prime}m^{-k+1}\|f_{j}\|_{W^{k,p}([a_{j}-2\epsilon,b_{j}+2\epsilon])}.$
Set, for example, $p_{m,j}=f_{j}(a_{j})+\int_{a_{j}}^{x}p_{m,j}^{\prime}(x^{\prime})\mathop{}\!\mathrm{d}x^{\prime}$. Then, define
@@
$\mathbf{Z}_{n}(x,y)=\mathbf{Y}_{n}(z;\mu)\begin{cases}\begin{bmatrix}1&0\\ -f_{j}^{\ddagger}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Omega_{j,+}^{\prime},\vskip 3.0pt\\ \vskip 3.0pt\begin{bmatrix}1&0\\ f_{j}^{\ddagger}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Omega_{j,-}^{\prime},\\ \mathbf{I}&\text{otherwise}.\end{cases}$
@@
$\overline{\partial}\mathbf{Z}_{n}(x,y)=\begin{cases}\mathbf{Z}_{n}(z;\mu)\begin{bmatrix}0&0\\ -\overline{\partial}r_{m,j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}&z\in\Omega_{j,+}^{\prime},\vskip 3.0pt\\ \vskip 3.0pt\mathbf{Z}_{n}(z;\mu)\begin{bmatrix}0&0\\ \overline{\partial}r_{m,j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}&z\in\Omega_{j,-}^{\prime},\\ \mathbf{0}&z\in\mathbb{C}\setminus\overline{\Omega}_{j}^{\prime}.\end{cases}$
@@
$\Sigma_{j,+}^{\prime}$ $=\{z~:~x=a_{j}-\epsilon/2,~0\leq y\leq\epsilon/2\}$
@@
i.e. the “top” boundary of the box $\Omega_{j}^{\prime}$. We give this negative orientation (as the boundary of $\Omega_{j}^{\prime})$. And set $\Sigma^{\prime}_{j,-}=\{z~:~\bar{z}\in\Sigma^{\prime}_{j,+}\}$, inheriting orientation.
@@
$\mathbf{Z}_{n}^{+}(x,y)=\mathbf{Z}_{n}^{-}(x,y)\begin{cases}\begin{bmatrix}1&0\\ f_{j}^{\ddagger}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Sigma_{j,-}^{\prime}\cup\Sigma_{j,+}^{\prime},\vskip 3.0pt\\ \begin{bmatrix}0&\rho(x)\\ -1/\rho(x)&0\end{bmatrix}&z\in[a_{j},b_{j}].\end{cases}$
@@
We find that $\mathbf{T}_{n}(x,y)$ solves the normalized hybrid RH-$\overline{\partial}$ problem
$\mathbf{T}_{n}^{+}(x,y)$ $=\mathbf{T}_{n}^{-}(x,y)\begin{cases}\begin{bmatrix}1&0\\ \,\mathrm{e}^{-2n\mathfrak{g}(z)}f_{j}^{\ddagger}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Sigma_{j,-}^{\prime}\cup\Sigma_{j,+}^{\prime},\vskip 3.0pt\\ \begin{bmatrix}0&\rho(x)\\ -1/\rho(x)&0\end{bmatrix}&z\in(a_{j},b_{j}),\vskip 3.0pt\\ \,\mathrm{e}^{-\Delta_{j}{\boldsymbol{\sigma}_{3}}}&z\in(b_{j},a_{j+1}),\end{cases}$
$\overline{\partial}\mathbf{T}_{n}(x,y)$ $=\mathbf{T}_{n}(z)\begin{cases}\begin{bmatrix}0&0\\ -\,\mathrm{e}^{-2n\mathfrak{g}(z)}\overline{\partial}r_{m,j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}&z\in\Omega_{j,+}^{\prime},\vskip 3.0pt\\ \vskip 3.0pt\begin{bmatrix}0&0\\ \,\mathrm{e}^{-2n\mathfrak{g}(z)}\overline{\partial}r_{m,j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}&z\in\Omega_{j,-}^{\prime},\\ \mathbf{0}&z\in\mathbb{C}\setminus\overline{\Omega_{j}^{\prime}},\end{cases}$
@@
$\mathbf{R}_{n}^{+}(x,y)=\mathbf{R}_{n}^{-}(x,y)\mathbf{G}_{n}(z)\begin{bmatrix}1&0\\ \,\mathrm{e}^{-2n\mathfrak{g}(z)}f_{j}^{\ddagger}(x,y)/\omega_{j}(z)&1\end{bmatrix}\mathbf{G}_{n}(z)^{-1},\quad z\in\Sigma_{j,-}^{\prime}\cup\Sigma_{j,+}^{\prime},$
$\overline{\partial}\mathbf{R}_{n}(x,y)$
$=\mathbf{R}_{n}(z)\begin{cases}\mathbf{G}_{n}(z)\,\mathrm{e}^{-{\boldsymbol{\sigma}_{3}}G(z)}\begin{bmatrix}0&0\\ -\,\mathrm{e}^{-2n\mathfrak{g}(z)-2G(z)}\overline{\partial}r_{m,j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}\,\mathrm{e}^{{\boldsymbol{\sigma}_{3}}G(z)}\mathbf{G}_{n}(z)^{-1}&z\in\Omega_{j,+}^{\prime},\vskip 3.0pt\\ \vskip 3.0pt\mathbf{G}_{n}(z)\,\mathrm{e}^{-{\boldsymbol{\sigma}_{3}}G(z)}\begin{bmatrix}0&0\\ \,\mathrm{e}^{-2n\mathfrak{g}(z)-2G(z)}\overline{\partial}r_{m,j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}\,\mathrm{e}^{{\boldsymbol{\sigma}_{3}}G(z)}\mathbf{G}_{n}(z)^{-1}&z\in\Omega_{j,-}^{\prime},\\ \mathbf{0}&z\in\mathbb{C}\setminus\overline{\Omega_{j}^{\prime}},\end{cases}$
@@
$\overline{\partial}\mathbf{C}_{n}(x,y)$ $=\mathbf{C}_{n}(z)\mathbf{X}_{n}(x,y),\quad z\in\Omega_{j}^{\prime}$
@@
Define the operator
$\mathcal{L}_{n}\mathbf{U}:=\mathcal{K}_{\Omega^{\prime}}(\mathbf{U}\mathbf{X}_{n}),$
and the companion operator
$\underline{\mathcal{K}}_{\Omega^{\prime}}u(z)=\frac{1}{\pi}\int_{\Omega^{\prime}}\frac{u(x^{\prime},y^{\prime})}{|z^{\prime}-z|}\mathop{}\!\mathrm{d}A(z^{\prime}).$
Below, we will establish conditions under which $\|\underline{\mathcal{K}}_{\Omega^{\prime}}(\|\mathbf{X}_{n}\|)\|_{L^{\infty}(\mathbb{C})}\to 0$. This will then imply, via a standard Neumann series argument, that $\mathbf{C}_{n}(x,y)$ and $\mathbf{C}_{n}(x,y)^{-1}$ are uniformly bounded and tend uniformly to the identity matrix.
Next, we make a similar observation to what was made in [40]. The Bernstein–Walsh inequality, see [32, Theorem III.2.1], in its simplest form, states that if a polynomial $p(x)$, of degree $n$, satisfies
@@
Thus, for the sequence of polynomials $p_{m,j}(x)$ that converge uniformly on $a_{j},b_{j}]$, we have $|p_{m,j}(x)|\leq C$ for some constant, independent of $m$ (and $j$). From this, and Proposition A.1, provided that $\epsilon$ is sufficiently small we have that
@@
For $k\geq 1$, $q>6$ and [^5] Recall that $\epsilon$ appears in the definition of $\Sigma_{j,\pm}^{\prime}$. $\epsilon$ sufficiently small set
$\mathbf{F}^{\natural}(x,y)$ $=\begin{cases}\begin{bmatrix}1&0\\ \pm f_{j}^{\ddagger}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Omega_{j,\pm}^{\prime},\\ \mathbf{I}&\text{otherwise},\end{cases}$
and suppose $\|\underline{\mathcal{K}}_{\Omega^{\prime}}(\|\mathbf{X}_{n}\|)\|_{L^{\infty}(\mathbb{C})}=o(1)$, $n\to\infty$. Then for $n$ sufficiently large,
@@
$\mathbf{V}_{n}(x,y)=\mathrm{O}\left(\frac{\|\underline{\mathcal{K}}_{\Omega^{\prime}}(\|\mathbf{X}_{n}\|)\|_{L^{\infty}(\mathbb{C})}}{1+|z|}\right),\quad z\in\mathbb{C},\quad m=\lfloor cn\rfloor.$
Proof.
The only remaining detail to spell out is the estimate on $\mathbf{V}_{n}$. First, it follows that
$\|\mathcal{L}_{n}\|_{L^{\infty}(\Omega^{\prime})}\leq\|\underline{\mathcal{K}}_{\Omega^{\prime}}(\|\mathbf{X}_{n}\|)\|_{L^{\infty}(\mathbb{C})}.$
This implies that the integral equation
$\mathbf{V}_{n}-\mathcal{L}_{n}\mathbf{V}_{n}=\mathcal{K}\mathbf{X}_{n}$
is near-identity. For $n$ sufficiently large, it has a unique solution, tending to zero as $n\to\infty$, and
$\|\mathbf{V}_{n}\|\leq\|\mathcal{L}_{n}\mathbf{V}_{n}\|+\|\underline{\mathcal{K}}_{\Omega^{\prime}}\mathbf{X}_{n}\|\leq(1+\|\mathbf{V}_{n}\|_{L^{\infty}(\Omega^{\prime})})\|\underline{\mathcal{K}}_{\Omega^{\prime}}(\|\mathbf{X}_{n}\|)\|_{L^{\infty}(\mathbb{C})}.$
Lemma 3.2 (Norm bounds).
For $p>2$, $q>6$, and $1/p=1/q+1/r$
$\|\underline{\mathcal{K}}_{\Omega^{\prime}}(\|\mathbf{X}_{n}\|)\|_{L^{\infty}(\mathbb{C})}=O(n^{-1/r}\max_{j}E_{j,m,q}).$
Furthermore,
$\|\underline{\mathcal{K}}_{\Omega^{\prime}}(\|\mathbf{X}_{n}\|)\|_{L^{\infty}(\mathbb{C})}=O(n^{-1}\log n\max_{j}E_{j,m,\infty}).$
Proof.
Importantly, we note that by Lemma C.3, $\mathrm{e}^{-2G(z)}/\omega_{j}(z)$ is bounded in a neighborhood of $a_{j},b_{j}]$. From the (at most) quarter-root singularities of $\mathbf{G}_{n}(z)\,\mathrm{e}^{-{\boldsymbol{\sigma}_{3}}G(z)}$ we find for $z\in\Omega_{j}^{\prime}$ that
$\mathbf{X}_{n}(x,y)=\mathbf{G}_{n}(z)\,\mathrm{e}^{-{\boldsymbol{\sigma}_{3}}G(z)}\begin{bmatrix}0&0\\ -\,\mathrm{e}^{-2n\mathfrak{g}(z)-2G(z)}\overline{\partial}r_{m,j}^{\sharp}(x,y)/\omega_{j}(z)&0\end{bmatrix}\,\mathrm{e}^{{\boldsymbol{\sigma}_{3}}G(z)}\mathbf{G}_{n}(z)^{-1}$
satisfies
$\|\mathbf{X}_{n}(x,y)\|\leq C|\overline{\partial}r_{m,j}^{\sharp}(x,y)||z-a_{j}|^{-1/2}|z-b_{j}|^{-1/2}\,\mathrm{e}^{-2n\operatorname{Re}\mathfrak{g}(z)}.$
With the aim of using (E.1), we have
(3.2) $\|\mathbf{X}_{n}\|_{L^{p}(\Omega_{j}^{\prime})}\leq C\|\overline{\partial}r_{m,j}^{\sharp}\|_{L^{q}(\Omega_{j}^{\prime})}\|h_{j,n}\|_{L^{r}(\Omega_{j}^{\prime})},\quad h_{j,n}(z):={|z-a_{j}|^{-1/2}}|z-b_{j}|^{-1/2}\,\mathrm{e}^{-2n\operatorname{Re}\mathfrak{g}(z)},$
for $1\leq p\leq 4$, and $1/p=1/q+1/r$. Since we will need $p>2$ below, and using (E.1) requires $\gamma<3/2$, we choose $r<3$, and $q>6$, giving
$\|\mathbf{X}_{n}\|_{L^{p}(\Omega_{j}^{\prime})}=O({n^{-1/r}}E_{j,m,q}).$
The first claim follows from the fact that $\|\underline{\mathcal{K}}_{\Omega^{\prime}}u\|_{L^{\infty}}(\mathbb{C})\leq C_{p}\|u\|_{L^{p}(\Omega^{\prime})}$ for some $C_{p}>0$, provided $p>2$.
To establish the second claim, we need to establish estimate
$\underline{\mathcal{K}}_{\Omega^{\prime}}h_{j,n}(z)=\mathrm{O}(n^{-1}\log n).$
This essentially follows from the proof of [40, Lemma 4] once we note that, after the change of variable $z^{\prime}=1/2(\tau+1/\tau)$, there is a helpful factor of $|\tau^{2}-1|$ that can be used to cancel the singularity in
$|z^{2}-1|^{-1/2}=\frac{2\tau}{|\tau^{2}-1|}.$